# Optical pumping of alkali-metal vapor with hyperfine-resolved buffer gas pressure

### 06 Mar 2026 v1

@@ IV.1 Absorption cross section
The results of $\sigma_{\mathrm{lin}}$ are shown in Fig. 5, where a notable feature emerges for finite $\Phi$. In the QHP regime, the resonance peaks of $\sigma_{\mathrm{lin}}$ for pumped atoms become significantly suppressed and unresolved. This behavior is evident in Fig. 5(a), while it is absent in the HP limit as shown in Fig. 5(b). A mechanism directly revealed by this is an effective pumping effect between F=a,bF=a,\,b multiplets. That is, when $\Delta_{F}=0$, population in the $F$ multiplet is more efficiently pumped into the other.
As another observable feature, in contrast to $\sigma_{0}$ in weak-pumping limit $\Phi\to 0$, we find that $\sigma_{\mathrm{lin}}$ in strong pumping limit $R_{\mathrm{op}}\gg\Gamma_{\mathrm{SD}}+\Gamma_{\mathrm{ex}}$ (but assure no population trapped in the ES) does not vanish but instead asymptotically approaches a finite value
$\sigma_{\mathrm{lin}}(R_{\mathrm{op}}\gg\Gamma_{\mathrm{SD}}+\Gamma_{\mathrm{ex}})$ $=\frac{A_{0}\Gamma_{\mathrm{brd}}}{(\Delta_{0}-\omega_{\mathrm{eff}})^{2}+\Gamma^{2}_{\mathrm{eff}}},$ (59)
which takes a Lorentzian form with $\omega_{\mathrm{eff}}$ and $\Gamma_{\mathrm{eff}}$ being the effective resonance frequency and linewidth,
$\omega_{\mathrm{eff}}$ $=\frac{[a]}{2[I]}\omega^{\{\mathrm{g}\}}_{a,0}+\frac{[b]}{2[I]}\omega^{\{\mathrm{g}\}}_{b,0}$, (60)
$\Gamma_{\mathrm{eff}}$ $=\Gamma_{\mathrm{brd}}^{2}+\frac{[a][b]}{2[I]}\left(\omega^{\{\mathrm{g}\}}_{a,0}+\omega^{\{\mathrm{g}\}}_{b,0}\right)^{2}.$ (61)
This behavior accounts for the convergence of the spectral area shown in Fig. 5(c), a characteristic that enables experimental reconstruction of the atomic state from the observed absorption spectrum.
Figure 5: (a) Absorption cross section $\sigma_{\mathrm{lin}}$ ($s=0$) versus laser detuning $\Delta_{0}$ with $\Gamma_{\mathrm{brd}}=3~\mathrm{GHz}$. The blue, green, and red curves correspond to $\Phi=10^{12}$, $10^{16}$, and $10^{20}~\mathrm{cm^{-2}\,s^{-1}}$, respectively. (b) Same as (a), but for a typical HP case, $2\Gamma_{\mathrm{brd}}=20~\mathrm{GHz}$. (c) Integrated spectral area obtained by $\int\sigma_{\mathrm{lin}}\,d\Delta_{0}$, plotted for different $\Gamma_{\mathrm{brd}}$. Relaxation rates, $\Gamma_{\mathrm{ex}}$ and $\Gamma_{\mathrm{SD}}$, correspond to 87Rb with $\mathrm{N}_{2}$ buffer gas at $100\,^{\circ}\mathrm{C}$.
Table 1: Expressions for $p_{F}(I,\,P)$ and $\langle S_{z}\rangle_{F}(I,\,P)$ defined by Eqs. (42) and (43), with $p_{b}=1-p_{a}$ and $\left\langle S_{z}\right\rangle_{b}=P/2-\left\langle S_{z}\right\rangle_{a}$, These results are derived in the STD limit and used in Eqs. (69) and (62).
II $p_{a}(I,P)$ $\langle S_{z}\rangle_{a}(I,P)$
$\dfrac{1}{2}$ $\dfrac{3}{4}+\dfrac{P^{2}}{4}$ $\dfrac{P}{2}$
11 $\dfrac{2}{3}+\dfrac{4P^{2}}{3(3+P^{2})}$ $\dfrac{5P+P^{3}}{3(3+P^{2})}$
$\dfrac{3}{2}$ $\dfrac{5}{8}+\dfrac{5P^{2}+P^{4}}{8(1+P^{2})}$ $\dfrac{5P+3P^{3}}{8\left(1+P^{2}\right)}$
22 $\dfrac{3}{5}+\dfrac{4(5P^{2}+3P^{4})}{5(5+10P^{2}+P^{4})}$ $\dfrac{3P+10P^{3}+3P^{5}}{2(5+10P^{2}+P^{4})}$
$\dfrac{5}{2}$ $\dfrac{7}{12}+\dfrac{35P^{2}+42P^{4}+3P^{6}}{12\left(3+10P^{2}+3P^{4}\right)}$ $\dfrac{7P+14P^{3}+3P^{5}}{3\left(3+10P^{2}+3P^{4}\right)}$
Figure 6: (a) Absorption cross section $\sigma_{\mathrm{cir}}$ of with $s=1$ and $2\Gamma_{\mathrm{brd}}=3~\mathrm{GHz}$. Three green solid curves, rendered from dark to light, correspond to photon flux $\Phi=0$, $5\times 10^{13}$, and $2\times 10^{14}~\mathrm{cm^{-2}\,s^{-1}}$, respectively. Dotted lines indicate the two characteristic laser detunings. (b) $\sigma_{\mathrm{cir}}$ as a function of $\Phi$ at $\Delta_{a}=0$ and $\Delta_{b}=0$. The intersection point of them $\sigma_{\mathrm{cir}}(\Delta_{a}=0)=\sigma_{\mathrm{cir}}(\Delta_{b}=0)$ occurs with $\Phi_{\mathrm{eq}}=7.5\times 10^{13}~\mathrm{cm^{-2}\,s^{-1}}$ where the polarization is about $P=0.2$. (c–d) The same conditions as in (a–b) are repeated for a larger collisional broadening $2\Gamma_{\mathrm{brd}}=20~\mathrm{GHz}$, representative of the HP limit. Relaxation rates, $\Gamma_{\mathrm{ex}}$ and $\Gamma_{\mathrm{SD}}$, correspond to 87Rb with $\mathrm{N}_{2}$ buffer gas at $100\,^{\circ}\mathrm{C}$.
Furthermore, for the circularly polarized photon absorption cross section $\sigma_{\mathrm{cir}}$, we adopt the STD limit mentioned in Sec. III.3.2, and are able to formally write $\sigma_{\mathrm{cir}}$ as
$\sigma_{\mathrm{cir}}$ $=C^{(\mathrm{cir})}_{a}\frac{A_{0}\Gamma_{\mathrm{brd}}}{\Delta_{a}^{2}+\Gamma_{\mathrm{brd}}^{2}}+C^{(\mathrm{cir})}_{b}\frac{A_{0}\Gamma_{\mathrm{brd}}}{\Delta_{b}^{2}+\Gamma_{\mathrm{brd}}^{2}},$ (62)
with $C^{(\mathrm{cir})}_{F}=p_{F}(I,\,P)-{2s\langle S_{z}\rangle_{F}}(I,\,P)$ for $F=a$ and $b$, being a function of the nuclear spin $I$ and the polarization $P$. Using Eq. (48), expressions of $p_{F}(I,\,P)$ and $\langle S_{z}\rangle_{F}(I,\,P)$ are both derived and displayed in Table 1. For $I=3/2$, as an example, the coefficients in Eq. (62) are
$C^{(\mathrm{cir})}_{a=2}$ $=\frac{5}{8}-\frac{10P-5P^{2}+6P^{3}-P^{4}}{8(1+P^{2})},$ (63)
$C^{(\mathrm{cir})}_{b=1}$ $=\frac{3}{8}+\frac{2P-5P^{2}-2P^{3}-P^{4}}{8(1+P^{2})},$ (64)
Typical results of $\sigma_{\mathrm{cir}}$ obtained by Eq. (40) are shown in Fig. 6(a) to illustrate how polarized atoms respond to pumping light with different detunings.
To clarify the optimal conditions for optical pumping efficiency, we focus on the relative magnitude of the cross sections at optical resonance, $\sigma_{\mathrm{cir}}(\Delta_{a}=0)$ and $\sigma_{\mathrm{cir}}(\Delta_{b}=0)$. It is noted that a critical point arises when the two cross sections become exactly equal, which occurs at the photon flux $\Phi=\Phi_{\mathrm{eq}}$, as shown in Fig. 6(b). For $\Phi$ larger than $\Phi_{\mathrm{eq}}$, atoms always exhibit stronger absorption of photons with $\Delta_{b}=0$, and consequently attain a higher polarization $P$ (as seen in Fig. 7), with the opposite behavior holding for $\Phi<\Phi_{\mathrm{eq}}$. This can be understood via how differently polarized atoms respond to the pumping mechanism. For low polarization, the higher degeneracy of the $a$ multiplet leads to a stronger response to pumping light with $\Delta_{a}=0$. As polarization builds up, the population is gradually transferred to the stretched dark state $|F=a,\,m_{F}=a\rangle$, while the remaining population, especially in the $b$ multiplet which contains opposite polarization induced by spin exchange, still absorbs photons. This renders resonant pumping at $\Delta_{b}=0$ favorable for further polarization, until atoms are almost fully polarized and consequently weaken $\sigma_{\mathrm{cir}}$ to zero. These different response behaviors can also be explained by $\mathrm{Im}\left[\delta H^{\{\mathrm{g}\}}\right]$ in Eq. (21), whose diagonal elements directly reflect the absorption strength of the corresponding states.
By contrast, in the HP limit, all the cross sections mentioned above collapse to
$\sigma^{(\mathrm{HP})}$ $=\frac{A_{0}\Gamma_{\mathrm{brd}}}{\Delta_{0}^{2}+\Gamma_{\mathrm{brd}}^{2}}\left(1-sP\right),$ (65)
which apparently takes a simple form but loses validity in the QHP regime.
IV.2 Electron spin polarization
Figure 7: Two–dimensional color map of the electron spin polarization $P$ versus $\Delta_{0}$ and $\Phi$. The vertical dotted lines indicate $\Delta_{a}=0$ and $\Delta_{b}=0$. The result is obtained by Eq. (69) and well agree with that in the STD limit. Contour lines of $P$ are overlaid for clarity. All additional atomic and optical parameters are identical to those used in Fig. 6.
Motivated by applications in atomic magnetometers, we primarily study the polarization $P$ in the QHP regime. We execute $(F_{z}|=(I_{z}|+(S_{z}|$ to both side of Eq. (44) and get an equation of the average total angular momentum $\langle F_{z}\rangle$
@@ IV.2 Electron spin polarization
$\frac{d\langle F_{z}\rangle}{dt}=$ $R_{\mathrm{op}}\sum_{F}\mathcal{Q}_{F}\left(\frac{s}{2}p_{F}-\langle S_{z}\rangle_{F}\right)-\Gamma_{\mathrm{SD}}\langle S_{z}\rangle,$ (66)
Here, we utilized the identities in the population subspace
$\left(F_{z}\left|\left[A_{\mathrm{SD}}\right]_{0,0}\right.\right.=\big(S_{z}\big|,$ (67)
$\left(F_{z}\left|\left[A^{(z)}_{\mathrm{SE}}\right]_{0,0}\right.\right.=\big(1^{\{\mathrm{g}\}}\big|,$ (68)
which physically means that the S-damping collision conserves the nuclear spin momentum, while the spin exchange conserves the total angular momentum (instead of $\langle S_{z}\rangle$, as presented in Sec. III.3.2). Thus, the nonlinear term in Eq. (45) vanishes for $\langle F_{z}\rangle$ evolution.
@@
Focusing on the steady state, we set $d\langle F_{z}\rangle/dt=0$ and get the polarization
@@
$P=\frac{sR_{\mathrm{op}}({Q}_{a}p_{a}+{Q}_{b}p_{b})}{\displaystyle R_{\mathrm{op}}\Big(\left\langle S_{z}\right\rangle_{a}{Q}_{a}+\left\langle S_{z}\right\rangle_{b}{Q}_{b}\Big)+\Gamma_{\mathrm{SD}}},$ (69)
which significantly differs from that in the HP limit ($Q_{F}\to 1$)
$P^{(\mathrm{HP})}=\dfrac{sR_{\mathrm{op}}}{R_{\mathrm{op}}+\Gamma_{\mathrm{SD}}}$ (70)
and depend on the atomic state in general. For simplicity and physical relevance, we still consider the STD limit verified above. With the expressions in Table 1, Eq. (69) becomes a polynomial equation of $P$ that can be solved only with experimental parameters. The result given by this is shown in Fig. 7, which primarily provides a qualitative reference for the overall trends of optical pumping efficiency.
Figure 8: (a) Magnetic resonance linewidth $\Gamma_{2}$ (see Eq. 52) for different laser detunings and collisional broadening values. The shaded regions highlight the narrower linewidth that arises when the laser is resonant with the $b$ multiplet. (b)–(c) show the normalized $X$ and $Y$ components in Eqs. (54) and (55), respectively, where the red and blue lines correspond to the marked points of the same colors in (a). The magnetic resonance signal at $\Delta_{b}=0$ exhibits a narrower linewidth and a stronger response. In above results, $\Gamma_{\mathrm{ex}}$ and $\Gamma_{\mathrm{SD}}$ correspond to 87Rb with $\mathrm{N}_{2}$ buffer gas at $100\,^{\circ}\mathrm{C}$, and a weak field $B=5~\mu\mathrm{T}$ is applied.

### 14 Apr 2026 v2

@@ IV.1 Absorption cross section
where the $F$-multiplet populations $p_{a}$ and $p_{b}$ are obtained in Eq. (41). The results of $\sigma_{\mathrm{lin}}$ are shown in Fig. 5(a), where a notable feature emerges for larger $\Phi$. In the QHP regime, the resonance peaks of $\sigma_{\mathrm{lin}}$ for pumped atoms become significantly suppressed and unresolved. This is evident in Fig. 5(a), while it is absent in the HP limit as shown in Fig. 5(b). The behavior can be understood from the fact that, when pumping near $\Delta_{F}=0$, population in a given hyperfine manifold $F$ is more efficiently pumped into the other manifold $F^{\prime}\neq F$, which leads to a reduction of the absorption. As the photon flux $\Phi$ continues to increase, this $F$-pumping mechanism is further enhanced, eventually driving $p_{F}$ almost entirely into a single hyperfine multiplet. Consequently, in the strong-pumping limit $R_{\mathrm{op}}\gg\Gamma_{\mathrm{SD}}+\Gamma_{\mathrm{ex}}$ (assuming negligible population trapping in the ES), the cross section asymptotically approaches a finite functional form:
$\sigma_{\mathrm{lin}}(R_{\mathrm{op}}\gg\Gamma_{\mathrm{SD}}+\Gamma_{\mathrm{ex}})$ $=\frac{A_{\mathrm{eff}}}{\pi}\frac{\Gamma_{\mathrm{eff}}}{\Delta^{2}_{0}+\Gamma^{2}_{\mathrm{eff}}},$ (46)
which exhibits a single-Lorentzian profile centered at the reference frequency, i.e., $\Delta_{0}=x_{a}\omega^{\{\mathrm{g}\}}_{a,0}+x_{b}\omega^{\{\mathrm{g}\}}_{b,0}=0$. Here, the effective spectral area and linewidth are $A_{\mathrm{eff}}=A/\xi$ and $\Gamma_{\mathrm{eff}}=\xi\,\Gamma_{\mathrm{brd}}$, respectively. The ratio $\xi$ is
$\xi$ $=\sqrt{1+x_{a}x_{b}\left(\dfrac{\omega^{\{\mathrm{g}\}}_{\mathrm{hf}}}{\Gamma_{\mathrm{brd}}}\right)^{2}},$ (47)
[…]
Throughout the process of varying $\Phi$, a notable feature is the emergence of a unique stationary point at the centered detuning
[…]
where $\sigma_{\mathrm{lin}}$ remains independent of $\Phi$. This characteristic point arises because the light is equally detuned from both multiplets ($\Delta_{a}=\Delta_{b}$). Such equal absorption strength effectively switches off the $F$-pumping dynamics, forcing the cross section to remain fixed at $\bar{\sigma}_{\mathrm{QHP}}$ Eq. (26).
Figure 5: Absorption cross section $\sigma$ versus light detuning. (a)–(b) Results for linear polarized pumping ($s=0$), under typical QHP ($2\Gamma_{\mathrm{brd}}=3~\mathrm{GHz}$) and HP ($2\Gamma_{\mathrm{brd}}=20~\mathrm{GHz}$) conditions. (c)–(d) Analogous plots for circular polarized pumping ($s=\pm 1$) on the $D_{1}$ line. In each panel, the vertical dotted lines denote the main resonance centers. The solid curves, arranged from top to bottom, span a photon flux $\Phi$ range from 0 to $10^{18}~\mathrm{cm^{-2}\cdot s^{-1}}$ for $s=0$, and from 0 to $5\times 10^{14}~\mathrm{cm^{-2}\cdot s^{-1}}$ for $s=\pm 1$. The 87Rb atoms are considered for simulations, with a fixed $\Gamma_{\mathrm{ex}}=7.75~\mathrm{kHz}$. To account for the spin-destruction collision induced by buffer gas, $\Gamma_{\mathrm{SD}}$ is set to $0.05~\mathrm{kHz}$ and $0.27~\mathrm{kHz}$ for the QHP and HP cases.
For circularly polarized pumping ($s=\pm 1$), we focus on the commonly used $D_{1}$ transition ($K=1$). Using the STD approximation detailed in Sec. III.3.2, we substitute the expressions of $p_{F}$ and $\langle S_{z}\rangle_{F}$ from Table 1 into (34). Consequently, the absorption cross section $\sigma_{\mathrm{cir}}$ is expressed as
$\sigma_{\mathrm{cir}}=A\left[C_{a}(I,\,P)\mathcal{L}(\Delta_{a})+C_{b}(I,\,P)\mathcal{L}(\Delta_{b})\right].$ (49)
where $C_{F}$ becomes an explicit function of nuclear spin $I$ and the polarization $P$. Taking $I=3/2$ as an example, the coefficients in Eq. (34) are
$C_{a=2}$ $=\frac{5}{8}-\frac{10P-5P^{2}+6P^{3}-P^{4}}{8(1+P^{2})},$ (50)
$C_{b=1}$ $=\frac{3}{8}+\frac{2P-5P^{2}-2P^{3}-P^{4}}{8(1+P^{2})},$ (51)
Representative spectra under QHP and HP conditions are displayed in Figs. 5(c) and 5(d), where an overall reduction in absorption as $\Phi$ increases is common to the both cases, mainly driven by the population accumulated in the dark stretched state $|F=a,m_{F}=sa\rangle$. Notably, Figure 5(c) shows an observation featured in the QHP regime: the higher resonance peak shifts from $\Delta_{a}=0$ at low $\Phi$ to $\Delta_{b}=0$ at high $\Phi$. The threshold photon flux equalizing these two peaks is denoted by $\Phi_{\mathrm{eq}}$ (solid black line). This inversion originates from two competing mechanisms. At low $\Phi$ values, low-polarization absorption is governed by the higher degeneracy of the $F=a$ multiplet (more available $m_{F}$ levels). As polarization builds up, the dominant absorbing states becomes $|F,m_{F}=s|a-1|\rangle$, and the state for $F=b$ contains a larger fraction of the electronic spin component capable of absorbing photons (e.g., spin-down electrons for $s=1$). The critical point of equal-absorption at $\Phi_{\mathrm{eq}}$ occurs when these two competing effects are balanced at $\Delta_{F}=0$ and lead to $C_{a}=C_{b}$. The corresponding polarization behavior under this condition is analyzed in the following subsection.
Table 1: Analytic expressions for $p_{a}$ and $\langle S_{z}\rangle_{a}$ defined in Eqs. (35) & (36), These relations are derived under the STD approximation, as a function of nuclear spin $I$ and the spin polarization $P$. The $F=b$ counterparts follow as $p_{b}=1-p_{a}$ and $\langle S_{z}\rangle_{b}=P/2-\langle S_{z}\rangle_{a}$. The rightmost column lists the equal-polarization point $P_{\mathrm{eq}}$ for D1 line pumping, obtained by solving Eq. (57) with $K=|s|=1$. For $I>1/2$, these results exclude the trivial $P=1$ root.
II $p_{a}(I,P)$ $\langle S_{z}\rangle_{a}(I,P)$ $|P_{\mathrm{eq}}|$
$\dfrac{1}{2}$ $\dfrac{3}{4}+\dfrac{P^{2}}{4}$ $\dfrac{P}{2}$ 1
11 $\dfrac{2}{3}+\dfrac{4P^{2}}{3(3+P^{2})}$ $\dfrac{5P+P^{3}}{3(3+P^{2})}$ 0.394
$\dfrac{3}{2}$ $\dfrac{5}{8}+\dfrac{5P^{2}+P^{4}}{8(1+P^{2})}$ $\dfrac{5P+3P^{3}}{8\left(1+P^{2}\right)}$ 0.207
22 $\dfrac{3}{5}+\dfrac{4(5P^{2}+3P^{4})}{5(5+10P^{2}+P^{4})}$ $\dfrac{3P+10P^{3}+3P^{5}}{2(5+10P^{2}+P^{4})}$ 0.128
$\dfrac{5}{2}$ $\dfrac{7}{12}+\dfrac{35P^{2}+42P^{4}+3P^{6}}{12\left(3+10P^{2}+3P^{4}\right)}$ $\dfrac{7P+14P^{3}+3P^{5}}{3\left(3+10P^{2}+3P^{4}\right)}$ 0.087
Beyond the specific spectral profiles against light detuning, the total light-absorption strength can be characterized by the integrated spectral area
$\alpha=\int\sigma(\Delta_{0})\,d\Delta_{0}.$ (52)
Figure 6 compares the variation of $\alpha$ for different $s$ in both QHP and HP cases. For linearly polarized light (black curves), $\alpha$ decays from $A$ and plateaus at $A_{\mathrm{eff}}$ in the strong-pumping limit, consistent with Eq. (46). Notably, the QHP case exhibits a pronounced reduction in $\alpha$ (solid line), whereas the HP case remains nearly unchanged (dashed line). This behavior is dictated by the ratio $\xi$ defined in Eq. (47)]: the HP limit $\Gamma_{\mathrm{brd}}\gg\omega^{\{\mathrm{g}\}}_{\mathrm{hf}}$ keeps $\xi\approx 1$ and suppresses $F$-pumping, whereas the QHP regime $\Gamma_{\mathrm{brd}}\sim\omega^{\{\mathrm{g}\}}_{\mathrm{hf}}$ allows active $F$-pumping to deplete absorption ($\xi>1$). Conversely, circular pumping (red curves) drives $\alpha$ strictly to zero, as $m_{F}$-pumping perfectly polarizes atoms into the non-absorbing stretched state. This vanishing is faster under QHP conditions due to stronger resonant absorption.
Figure 6: Integrated spectral area $\alpha$ [see Eq. (52)] normalized by AA versus photon flux $\Phi$, incorporating both linear ($s=0$) and circular ($s=\pm 1$, $D_{1}$ line) polarized pumping. Solid and dashed lines correspond to the QHP and HP cases, respectively. With increasing $\Phi$, the $s=0$ curves asymptotically converge to $\xi=0.41$ and $0.95$ [see Eq. (47)]. Parameters are consistent with Fig. 5.
IV.2 Spin polarization
Building upon the population dynamics in Sec. III.3.2, we study the steady-state spin polarization $P$ in the QHP regime. To start with, applying $(F_{z}|=(I_{z}|+(S_{z}|$ to both sides of Eq. (37) yields the evolution equation for the average total angular momentum $\langle F_{z}\rangle$:
@@ IV.2 Spin polarization
$\frac{d\langle F_{z}\rangle}{dt}=$ $2R_{\mathrm{op}}\sum_{F}\mathcal{Q}_{F}\left(Ksp_{F}-2\langle S_{z}\rangle_{F}\right)-\Gamma_{\mathrm{SD}}\langle S_{z}\rangle,$ (53)
In obtaining this equation, we apply the superoperator identities in the population subspace
[…]
by which the nonlinear $\Gamma_{\mathrm{ex}}$ term vanishes in the $\langle F_{z}\rangle$ evolution, and meanwhile, the optical pumping superoperator $\left[A_{\mathrm{op}}\right]_{0,0}$ contributes the $p_{F}$ and $\langle S_{z}\rangle_{F}$ terms. They effectively acts as $F$-resolved spin-exchange and spin-destruction processes, which respectively capture the distinct impacts of $F$-pumping and $m_{F}$-pumping on the overall polarization dynamics.
Figure 7: (a) Two–dimensional color map of spin polarization $P$, with $K=|s|=1$ and $2\Gamma_{\mathrm{brd}}=3$ GHz. The white line signifies the photon flux $\Phi^{\prime}_{\mathrm{eq}}=7.6\times 10^{12}~\mathrm{cm^{-2}\cdot s^{-1}}$ that leads to $P(\Delta_{a}=0)=P(\Delta_{b}=0)=P_{\mathrm{eq}}$. (b) Variation of polarization with light detuning at representative $\Phi_{\mathrm{}}$ values. The black dashed line isolates the $\Phi^{\prime}_{\mathrm{eq}}$ case, alongside the solid red and blue curves for $\Phi=2\times 10^{12}$ and $2\times 10^{13}~\mathrm{cm^{-2}\cdot s^{-1}}$. The spin-destruction rate for 87Rb atoms used here is fixed at $\Gamma_{\mathrm{SD}}=0.05~\mathrm{kHz}$.
@@
By setting $d\langle F_{z}\rangle/dt=0$, we obtain the implicit steady-state equation for the polarization:
$P=\frac{KsR_{\mathrm{op}}({Q}_{a}p_{a}+{Q}_{b}p_{b})}{\displaystyle R_{\mathrm{op}}\left({Q}_{a}u_{a}+{Q}_{b}u_{b}\right)+\Gamma_{\mathrm{SD}}},$ (56)
where $u_{F}=\left\langle S_{z}\right\rangle_{F}/\left\langle S_{z}\right\rangle$ represents the polarization weight of the $F$ multiplet. Adopting the STD approximation, we utilize the explicit expressions in Table 1 to convert Eq. (56) into a polynomial equation of $P$, which can be readily solved once the rate parameters are given. Under typical QHP conditions, the resulting $P$ is shown in Fig. 7, revealing several distinct features. Figure 7(a) demonstrates that $P$ generally builds up with increasing $\Phi$, and manifests two distinct maxima at $\Delta_{a}=0$ and $\Delta_{b}=0$ when the hyperfine structure are well-resolved. Figure 7(b) profiles several representative slices, highlights a shifting dominance across distinct $\Phi$ levels: lower photon fluxes favor polarization via the $\Delta_{a}=0$ resonance (red line), whereas higher fluxes make $\Delta_{b}=0$ more efficient (blue line). This behavior mirrors the $D_{1}$ absorption cross section $\sigma_{\mathrm{cir}}$ disscussed in Sec. IV.1; stronger resonant absorption directly dictates a higher rate of angular momentum injection, thus yielding a higher $P$.
Importantly, an equal-polarization point emerges at a critical flux $\Phi=\Phi^{\prime}_{\mathrm{eq}}$, where the polarization values at the $F=a$ and $F=b$ resonances equalize at a common value $P_{\mathrm{eq}}$. This phenomenon is denoted by the white dashed line in Fig. 7(a) and the black curve in Fig. 7(b). From Eq. (56), solving for the $P_{\mathrm{eq}}$ critical condition is equivalent to
$Ksp_{a}-2\langle S_{z}\rangle_{a}$ $=Ksp_{b}-2\langle S_{z}\rangle_{b},$ (57)
with the corresponding flux calculated as
@@
$\Phi^{\prime}_{\mathrm{eq}}$ $=\frac{P_{\mathrm{eq}}\Gamma_{\mathrm{SD}}}{A(Ksp_{F}-2\langle S_{z}\rangle_{F})[\mathcal{L}(0)+\mathcal{L}(\omega^{\{\mathrm{g}\}}_{\mathrm{hf}})]}$. (58)
Here, Equation (57) signifies a fundamental balance independent of collisional broadening or any relaxation rate. This physically indicates that the combined optical pumping effects (including $F$- and $m_{F}$-pumping) exert an identical impact on $P$ through $F=a$ and $F=b$ levels. Under the STD approximation, Eq. (57) becomes a polynomial equation of $P_{\mathrm{eq}}$ uniquely determined by the nuclear spin $I$ (as listed in Table 1). We observe that a larger $I$ leads to a higher $P_{\mathrm{eq}}$. This stems from the fact that a larger $I$ creates more $m_{F}$ sublevels, which dilutes the initial low-$\Phi$ degeneracy advantage of $F=a$, allowing the $F=b$ resonance to rapidly overtake it in polarization efficiency.
Finally, it is worth noting that for the $D_{1}$ line ($K=1$), this equal-polarization state is exactly identical to the equal-absorption point $(C_{a}=C_{b})$, obtained by the common photon flux $\Phi_{\mathrm{eq}}=\Phi^{\prime}_{\mathrm{eq}}$. However, this correspondence does not hold for the $D_{2}$ line (K=1/2K=-1/2) due to its fundamentally different optical pumping pathways.
Figure 8: (a) Light-induced contribution to the magnetic resonance linewidth, $\delta\Gamma_{2}$, evaluated at D1 line optical resonance. The solid and dashed curves denote the QHP and HP cases. The shaded regions highlight the narrowing effect when the pumping light targets the $F=b$ level in the QHP regime. (b) and (c) display the normalized in-phase and out-of-phase quadratures ($X$ and $Y$) of the complex response $R$ [see Eq. (44)]. The red and blue curves correspond to the colored markers in (a), representing $\Delta_{a}=0$ and $\Delta_{b}=0$. For these calculations, we consider 87Rb atoms subjected to a weak magnetic field $B=5~\mu\mathrm{T}$, with detailed relaxation parameters matching those in figures in Sec. V.